A 2020 kg car traveling to the west at 22.2
m/s slows down uniformly.

a) How long would it take the car to come to
a stop if the force on the car is 8840 N to the
east? Let East be positive.
Answer in units of s.

b) What is the car’s displacement during the
time it takes to stop?
Answer in units of m.
1
Expert's answer
2020-04-13T10:21:52-0400

a)


"t=\frac{v}{a}=\frac{mv}{F}=\frac{(2020)(22.2)}{8840}=5.07\ s"

b)


"d=-\frac{v^2}{2a}=-\frac{(2020)(22.2)^2}{2\cdot8840}=-56.3\ m"

Negative sign means that the direction of displacement is westwards.